# Chain-linked Multiple Matrix Integration via Embedding Alignment

### v2

@@ Appendix D Technical Lemmas
For $\hat{\bm{\Lambda}}^{1/2}\mathbf{W}_{\mathbf{U}}-\mathbf{W}_{\mathbf{U}}\bm{\Lambda}^{1/2}$, for any $k,l\in[d]$ the $(k,l)$ entry can be written as
$(\hat{\bm{\Lambda}}^{1/2}\mathbf{W}_{\mathbf{U}}-\mathbf{W}_{\mathbf{U}}\bm{\Lambda}^{1/2})_{k,l}$ $=(\mathbf{W}_{\mathbf{U}})_{k,l}\cdot\big{(}\sqrt{\lambda_{l}(\mathbf{A})}-\sqrt{\lambda_{k}(\mathbf{P})}\big{)}$ (D.11)
@@
$=(\hat{\bm{\Lambda}}\mathbf{W}_{\mathbf{U}}-\mathbf{W}_{\mathbf{U}}\bm{\Lambda})_{k,l}\cdot\big{(}\sqrt{\lambda_{l}(\mathbf{A})}+\sqrt{\lambda_{k}(\mathbf{P})}\big{)}^{-1}.$
We define $\mathbf{H}$ as a $d\times d$ matrix whose entries are $\mathbf{H}_{k,\ell}=\big{(}\sqrt{\lambda_{\ell}(\mathbf{A})}+\sqrt{\lambda_{k}(\mathbf{P})}\big{)}^{-1}$. Eq. (D.11) means
$\hat{\bm{\Lambda}}^{1/2}\mathbf{W}_{\mathbf{U}}-\mathbf{W}_{\mathbf{U}}\bm{\Lambda}^{1/2}=(\hat{\bm{\Lambda}}\mathbf{W}_{\mathbf{U}}-\mathbf{W}_{\mathbf{U}}\bm{\Lambda})\circ\mathbf{H},$
where $\circ$ denotes the Hadamard matrix product, and by Eq. (D.10) and Lemma D.2 it follows that
$\|\hat{\bm{\Lambda}}^{1/2}\mathbf{W}_{\mathbf{U}}-\mathbf{W}_{\mathbf{U}}\bm{\Lambda}^{1/2}\|$ $\leq d^{1/2}\|\mathbf{H}\|_{\max}\cdot\|\hat{\bm{\Lambda}}\mathbf{W}_{\mathbf{U}}-\mathbf{W}_{\mathbf{U}}\bm{\Lambda}\|$ (D.12)
$\leq d^{1/2}(\sqrt{\lambda_{d}(\mathbf{A})}+\sqrt{\lambda_{d}(\mathbf{P})})^{-1}\cdot\|\bm{\Lambda}\mathbf{W}_{\mathbf{U}}^{\top}-\mathbf{W}_{\mathbf{U}}^{\top}\hat{\bm{\Lambda}}\|$
@@
$\breve{\mathbf{W}}=\underset{\mathbf{O}\in\mathbb{O}_{d}}{\arg\min}\|\hat{\mathbf{X}}\mathbf{O}-\mathbf{X}\|_{F}\text{ and }\mathbf{W}_{\mathbf{U}}=\hat{H}(\mathbf{W}_{\mathbf{U}}\bm{\Lambda}),$
where $\hat{H}(\cdot)$ is a matrix-valued function, and for any $d\times d$ invertible matrix $\mathbf{C}$, $\hat{H}(\mathbf{C})=\mathbf{C}(\mathbf{C}^{\top}\mathbf{C})^{-1/2}$. Then if $\|\hat{\mathbf{X}}^{\top}\mathbf{X}-\mathbf{W}_{\mathbf{U}}\bm{\Lambda}\|\leq\lambda_{d}(\mathbf{W}_{\mathbf{U}}\bm{\Lambda})$, according to Theorem 1 in Li (1995) we have
$\|\breve{\mathbf{W}}-\mathbf{W}_{\mathbf{U}}\|$ $\leq\frac{2\|\hat{\mathbf{X}}^{\top}\mathbf{X}-\mathbf{W}_{\mathbf{U}}\bm{\Lambda}\|}{\sigma_{d}(\hat{\mathbf{X}}^{\top}\mathbf{X})+\sigma_{d}(\mathbf{W}_{\mathbf{U}}\bm{\Lambda})}\leq\frac{2\|\hat{\mathbf{X}}^{\top}\mathbf{X}-\mathbf{W}_{\mathbf{U}}\bm{\Lambda}\|}{\lambda_{i,\min}}$
We now bound $\|\hat{\mathbf{X}}^{\top}\mathbf{X}-\mathbf{W}_{\mathbf{U}}\bm{\Lambda}\|$. By Lemma D.2 and Eq. (D.12) we have
$\|\hat{\mathbf{X}}^{\top}\mathbf{X}-\mathbf{W}_{\mathbf{U}}\bm{\Lambda}\|$ $=\|\hat{\bm{\Lambda}}^{1/2}\hat{\mathbf{U}}^{\top}\mathbf{U}\bm{\Lambda}^{1/2}-\mathbf{W}_{\mathbf{U}}\bm{\Lambda}\|$
$=\|\hat{\bm{\Lambda}}^{1/2}(\hat{\mathbf{U}}^{\top}\mathbf{U}-\mathbf{W}_{\mathbf{U}})\bm{\Lambda}^{1/2}+(\hat{\bm{\Lambda}}^{1/2}\mathbf{W}_{\mathbf{U}}-\mathbf{W}_{\mathbf{U}}\bm{\Lambda}^{1/2})\bm{\Lambda}^{1/2}\|$
$\leq\|\hat{\bm{\Lambda}}\|^{1/2}\cdot\|\hat{\mathbf{U}}^{\top}\mathbf{U}-\mathbf{W}_{\mathbf{U}}\|\cdot\|\bm{\Lambda}\|^{1/2}+\|\hat{\bm{\Lambda}}^{1/2}\mathbf{W}_{\mathbf{U}}-\mathbf{W}_{\mathbf{U}}\bm{\Lambda}^{1/2}\|\cdot\|\bm{\Lambda}\|^{1/2}$
$\lesssim\lambda_{i,\max}\cdot\frac{(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})^{2}n_{i}}{q_{i}\lambda_{i,\min}^{2}}$
$+\Big{(}\frac{d^{1/2}(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})^{2}n_{i}}{q_{i}\lambda_{i,\min}^{3/2}}+\frac{d(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})\log^{1/2}n_{i}}{q_{i}^{1/2}\lambda_{i,\min}^{1/2}}\Big{)}\cdot\lambda_{i,\max}^{1/2}$
@@
with high probability. Now by Eq. (C.1) we have $\|\hat{\mathbf{X}}^{\top}\mathbf{X}-\mathbf{W}_{\mathbf{U}}\bm{\Lambda}\|\leq\lambda_{i,\min}$ with high probability. In summary we have
$\|\breve{\mathbf{W}}-\mathbf{W}_{\mathbf{U}}\|$ $\leq\frac{2\|\hat{\mathbf{X}}^{\top}\mathbf{X}-\mathbf{W}_{\mathbf{U}}\bm{\Lambda}\|}{\lambda_{i,\min}}\lesssim\frac{d^{1/2}(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})^{2}n_{i}}{q_{i}\lambda_{i,\min}^{2}}+\frac{d(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})\log^{1/2}n_{i}}{q_{i}^{1/2}\lambda_{i,\min}}$
@@
D.2 Proof of Lemma C.1
We first state a lemma for bounding the error of $\hat{\mathbf{U}}^{(i)}$ as an estimate of $\mathbf{U}^{(i)}$; see Section D.3 for a proof.
Lemma D.4.
Consider the setting of Lemma C.1. Define $\mathbf{W}^{(i)}_{\mathbf{U}}$ as a minimizer of $\|\hat{\mathbf{U}}^{(i)}\mathbf{O}-\mathbf{U}^{(i)}\|_{F}$ over all $d\times d$ orthogonal matrices $\mathbf{O}$. We then have
$\hat{\mathbf{U}}^{(i)}\mathbf{W}^{(i)}_{\mathbf{U}}-\mathbf{U}^{(i)}=\mathbf{E}^{(i)}\mathbf{U}^{(i)}(\bm{\Lambda}^{(i)})^{-1}+\mathbf{R}_{\mathbf{U}}^{(i)},$ (D.13)
@@
$\mathbf{X}_{\mathcal{U}_{i}}=\mathbf{U}_{\mathcal{U}_{i}}\bm{\Lambda}^{1/2},\quad\mathbf{X}^{(i)}=\mathbf{U}^{(i)}(\bm{\Lambda}^{(i)})^{1/2}\text{ and }\hat{\mathbf{X}}^{(i)}=\hat{\mathbf{U}}^{(i)}(\hat{\bm{\Lambda}}^{(i)})^{1/2}.$
As $\mathbf{X}_{\mathcal{U}_{i}}\mathbf{X}_{\mathcal{U}_{i}}^{\top}=\mathbf{P}^{(i)}=\mathbf{X}^{(i)}\mathbf{X}^{(i)\top}$, there exists an orthogonal $\tilde{\mathbf{W}}^{(i)}$ such that $\mathbf{X}_{\mathcal{U}_{i}}=\mathbf{X}^{(i)}\tilde{\mathbf{W}}^{(i)}$. Define
$\breve{\mathbf{W}}^{(i)}=\underset{\mathbf{O}\in\mathbb{O}_{d}}{\arg\min}\|\hat{\mathbf{X}}^{(i)}\mathbf{O}-\mathbf{X}^{(i)}\|_{F},$ (D.14)
@@
$\mathbf{W}^{(i)}=\underset{\mathbf{O}\in\mathbb{O}_{d}}{\arg\min}\|\hat{\mathbf{X}}^{(i)}\mathbf{O}-\mathbf{X}_{\mathcal{U}_{i}}\|_{F}\text{ and }\mathbf{W}^{(i)}_{\mathbf{U}}=\underset{\mathbf{O}\in\mathbb{O}_{d}}{\arg\min}\|\hat{\mathbf{U}}^{(i)}\mathbf{O}-\mathbf{U}^{(i)}\|_{F}.$
Note that $\mathbf{W}^{(i)}=\breve{\mathbf{W}}^{(i)}\tilde{\mathbf{W}}^{(i)}$. Next, by Eq. (D.13) we have
$\hat{\mathbf{X}}^{(i)}\mathbf{W}^{(i)}-\mathbf{X}_{\mathcal{U}_{i}}=$ $[\hat{\mathbf{X}}^{(i)}\breve{\mathbf{W}}^{(i)}-\mathbf{X}^{(i)}]\tilde{\mathbf{W}}^{(i)}$
$=$ $\Big{[}[(\hat{\mathbf{U}}^{(i)}\mathbf{W}_{\mathbf{U}}^{(i)}-\mathbf{U}^{(i)})+\mathbf{U}^{(i)}](\mathbf{W}_{\mathbf{U}}^{(i)})^{\top}(\hat{\bm{\Lambda}}^{(i)})^{1/2}(\breve{\mathbf{W}}^{(i)}-\mathbf{W}^{(i)}_{\mathbf{U}})$
$+[(\hat{\mathbf{U}}^{(i)}\mathbf{W}_{\mathbf{U}}^{(i)}-\mathbf{U}^{(i)})+\mathbf{U}^{(i)}](\mathbf{W}_{\mathbf{U}}^{(i)})^{\top}[(\hat{\bm{\Lambda}}^{(i)})^{1/2}\mathbf{W}^{(i)}_{\mathbf{U}}-\mathbf{W}^{(i)}_{\mathbf{U}}(\bm{\Lambda}^{(i)})^{1/2}]$
$+(\hat{\mathbf{U}}^{(i)}\mathbf{W}^{(i)}_{\mathbf{U}}-\mathbf{U}^{(i)})(\bm{\Lambda}^{(i)})^{1/2}\Big{]}\tilde{\mathbf{W}}^{(i)}$
$=$ $\mathbf{E}^{(i)}\mathbf{U}^{(i)}(\bm{\Lambda}^{(i)})^{-1/2}\tilde{\mathbf{W}}^{(i)}+\mathbf{R}^{(i)}$
$=$ $\mathbf{E}^{(i)}\mathbf{X}^{(i)}(\mathbf{X}^{(i)\top}\mathbf{X}^{(i)})^{-1}\tilde{\mathbf{W}}^{(i)}+\mathbf{R}^{(i)}$
@@
$\mathbf{R}^{(i)}=$ $\Big{[}\mathbf{R}^{(i)}_{\mathbf{U}}(\bm{\Lambda}^{(i)})^{1/2}+[(\hat{\mathbf{U}}^{(i)}\mathbf{W}_{\mathbf{U}}^{(i)}-\mathbf{U}^{(i)})+\mathbf{U}^{(i)}](\mathbf{W}_{\mathbf{U}}^{(i)})^{\top}(\hat{\bm{\Lambda}}^{(i)})^{1/2}(\breve{\mathbf{W}}^{(i)}-\mathbf{W}^{(i)}_{\mathbf{U}})$
$+[(\hat{\mathbf{U}}^{(i)}\mathbf{W}_{\mathbf{U}}^{(i)}-\mathbf{U}^{(i)})+\mathbf{U}^{(i)}](\mathbf{W}_{\mathbf{U}}^{(i)})^{\top}[(\hat{\bm{\Lambda}}^{(i)})^{1/2}\mathbf{W}^{(i)}_{\mathbf{U}}-\mathbf{W}^{(i)}_{\mathbf{U}}(\bm{\Lambda}^{(i)})^{1/2}]\Big{]}\tilde{\mathbf{W}}^{(i)}.$
By Lemma D.4, Lemma D.2 and Lemma D.3 we obtain
$\|\mathbf{R}^{(i)}\|$ $\lesssim\|\mathbf{R}^{(i)}_{\mathbf{U}}\|\cdot\|\bm{\Lambda}^{(i)}\|^{1/2}+[\|\hat{\mathbf{U}}^{(i)}\mathbf{W}_{\mathbf{U}}^{(i)}-\mathbf{U}^{(i)}\|+\|\mathbf{U}^{(i)}\|]\cdot\|\hat{\bm{\Lambda}}^{(i)}\|^{1/2}\cdot\|\breve{\mathbf{W}}^{(i)}-\mathbf{W}^{(i)}_{\mathbf{U}}\|$
$+[\|\hat{\mathbf{U}}^{(i)}\mathbf{W}_{\mathbf{U}}^{(i)}-\mathbf{U}^{(i)}\|+\|\mathbf{U}^{(i)}\|]\cdot\|(\hat{\bm{\Lambda}}^{(i)})^{1/2}\mathbf{W}^{(i)}_{\mathbf{U}}-\mathbf{W}^{(i)}_{\mathbf{U}}(\bm{\Lambda}^{(i)})^{1/2}\|$
@@
$\|\mathbf{R}^{(i)}\|_{2\to\infty}$ $\lesssim\|\mathbf{R}^{(i)}_{\mathbf{U}}\|_{2\to\infty}\cdot\|\bm{\Lambda}^{(i)}\|^{1/2}+[\|\hat{\mathbf{U}}^{(i)}\mathbf{W}_{\mathbf{U}}^{(i)}-\mathbf{U}^{(i)}\|_{2\to\infty}+\|\mathbf{U}^{(i)}\|_{2\to\infty}]\cdot\|\hat{\bm{\Lambda}}^{(i)}\|^{1/2}\cdot\|\breve{\mathbf{W}}^{(i)}-\mathbf{W}^{(i)}_{\mathbf{U}}\|$
$+[\|\hat{\mathbf{U}}^{(i)}\mathbf{W}_{\mathbf{U}}^{(i)}-\mathbf{U}^{(i)}\|_{2\to\infty}+\|\mathbf{U}^{(i)}\|_{2\to\infty}]\cdot\|(\hat{\bm{\Lambda}}^{(i)})^{1/2}\mathbf{W}^{(i)}_{\mathbf{U}}-\mathbf{W}^{(i)}_{\mathbf{U}}(\bm{\Lambda}^{(i)})^{1/2}\|$
@@
$\hat{\mathbf{U}}$ $=\mathbf{A}\hat{\mathbf{U}}\hat{\bm{\Lambda}}^{-1}=(\mathbf{U}\bm{\Lambda}\mathbf{U}^{\top}+\mathbf{E})\hat{\mathbf{U}}\hat{\bm{\Lambda}}^{-1}=\mathbf{U}\mathbf{U}^{\top}\hat{\mathbf{U}}+\mathbf{U}(\bm{\Lambda}\mathbf{U}^{\top}\hat{\mathbf{U}}-\mathbf{U}^{\top}\hat{\mathbf{U}}\hat{\bm{\Lambda}})\hat{\bm{\Lambda}}^{-1}+\mathbf{E}\hat{\mathbf{U}}\hat{\bm{\Lambda}}^{-1}.$

### v4

@@ Appendix E Technical Lemmas
For $\widehat{\bm{\Lambda}}^{1/2}\mathbf{W}_{\mathbf{U}}-\mathbf{W}_{\mathbf{U}}\bm{\Lambda}^{1/2}$, for any $k,l\in[d]$ the $(k,l)$ entry can be written as
$(\widehat{\bm{\Lambda}}^{1/2}\mathbf{W}_{\mathbf{U}}-\mathbf{W}_{\mathbf{U}}\bm{\Lambda}^{1/2})_{k,l}$ $=(\mathbf{W}_{\mathbf{U}})_{k,l}\cdot\big{(}\sqrt{\lambda_{l}(\mathbf{A})}-\sqrt{\lambda_{k}(\mathbf{P})}\big{)}$ (E.11)
@@
$=(\widehat{\bm{\Lambda}}\mathbf{W}_{\mathbf{U}}-\mathbf{W}_{\mathbf{U}}\bm{\Lambda})_{k,l}\cdot\big{(}\sqrt{\lambda_{l}(\mathbf{A})}+\sqrt{\lambda_{k}(\mathbf{P})}\big{)}^{-1}.$
We define $\mathbf{H}$ as a $d\times d$ matrix whose entries are $\mathbf{H}_{k,\ell}=\big{(}\sqrt{\lambda_{\ell}(\mathbf{A})}+\sqrt{\lambda_{k}(\mathbf{P})}\big{)}^{-1}$. Eq. (E.11) means
$\widehat{\bm{\Lambda}}^{1/2}\mathbf{W}_{\mathbf{U}}-\mathbf{W}_{\mathbf{U}}\bm{\Lambda}^{1/2}=(\widehat{\bm{\Lambda}}\mathbf{W}_{\mathbf{U}}-\mathbf{W}_{\mathbf{U}}\bm{\Lambda})\circ\mathbf{H},$
where $\circ$ denotes the Hadamard matrix product, and by Eq. (E.10) and Lemma E.2 it follows that
$\|\widehat{\bm{\Lambda}}^{1/2}\mathbf{W}_{\mathbf{U}}-\mathbf{W}_{\mathbf{U}}\bm{\Lambda}^{1/2}\|$ $\leq d^{1/2}\|\mathbf{H}\|_{\max}\cdot\|\widehat{\bm{\Lambda}}\mathbf{W}_{\mathbf{U}}-\mathbf{W}_{\mathbf{U}}\bm{\Lambda}\|$ (E.12)
$\leq d^{1/2}(\sqrt{\lambda_{d}(\mathbf{A})}+\sqrt{\lambda_{d}(\mathbf{P})})^{-1}\cdot\|\bm{\Lambda}\mathbf{W}_{\mathbf{U}}^{\top}-\mathbf{W}_{\mathbf{U}}^{\top}\widehat{\bm{\Lambda}}\|$
@@
$\breve{\mathbf{W}}=\underset{\mathbf{O}\in\mathcal{O}_{d}}{\arg\min}\|\widehat{\mathbf{X}}\mathbf{O}-\mathbf{X}\|_{F}\text{ and }\mathbf{W}_{\mathbf{U}}=\widehat{H}(\mathbf{W}_{\mathbf{U}}\bm{\Lambda}),$
[…]
$\|\breve{\mathbf{W}}-\mathbf{W}_{\mathbf{U}}\|$ $\leq\frac{2\|\widehat{\mathbf{X}}^{\top}\mathbf{X}-\mathbf{W}_{\mathbf{U}}\bm{\Lambda}\|}{\sigma_{d}(\widehat{\mathbf{X}}^{\top}\mathbf{X})+\sigma_{d}(\mathbf{W}_{\mathbf{U}}\bm{\Lambda})}\leq\frac{2\|\widehat{\mathbf{X}}^{\top}\mathbf{X}-\mathbf{W}_{\mathbf{U}}\bm{\Lambda}\|}{\lambda_{i,\min}}$
We now bound $\|\widehat{\mathbf{X}}^{\top}\mathbf{X}-\mathbf{W}_{\mathbf{U}}\bm{\Lambda}\|$. By Lemma E.2 and Eq. (E.12) we have
$\|\widehat{\mathbf{X}}^{\top}\mathbf{X}-\mathbf{W}_{\mathbf{U}}\bm{\Lambda}\|$ $=\|\widehat{\bm{\Lambda}}^{1/2}\widehat{\mathbf{U}}^{\top}\mathbf{U}\bm{\Lambda}^{1/2}-\mathbf{W}_{\mathbf{U}}\bm{\Lambda}\|$
$=\|\widehat{\bm{\Lambda}}^{1/2}(\widehat{\mathbf{U}}^{\top}\mathbf{U}-\mathbf{W}_{\mathbf{U}})\bm{\Lambda}^{1/2}+(\widehat{\bm{\Lambda}}^{1/2}\mathbf{W}_{\mathbf{U}}-\mathbf{W}_{\mathbf{U}}\bm{\Lambda}^{1/2})\bm{\Lambda}^{1/2}\|$
$\leq\|\widehat{\bm{\Lambda}}\|^{1/2}\cdot\|\widehat{\mathbf{U}}^{\top}\mathbf{U}-\mathbf{W}_{\mathbf{U}}\|\cdot\|\bm{\Lambda}\|^{1/2}+\|\widehat{\bm{\Lambda}}^{1/2}\mathbf{W}_{\mathbf{U}}-\mathbf{W}_{\mathbf{U}}\bm{\Lambda}^{1/2}\|\cdot\|\bm{\Lambda}\|^{1/2}$
$\lesssim\lambda_{i,\max}\cdot\frac{(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})^{2}n_{i}}{q_{i}\lambda_{i,\min}^{2}}+\Big{(}\frac{d^{1/2}(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})^{2}n_{i}}{q_{i}\lambda_{i,\min}^{3/2}}+\frac{d(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})\log^{1/2}n_{i}}{q_{i}^{1/2}\lambda_{i,\min}^{1/2}}\Big{)}\cdot\lambda_{i,\max}^{1/2}$
@@
with high probability. Now by Eq. (D.1) we have $\|\widehat{\mathbf{X}}^{\top}\mathbf{X}-\mathbf{W}_{\mathbf{U}}\bm{\Lambda}\|\leq\lambda_{i,\min}$ with high probability. In summary we have
$\|\breve{\mathbf{W}}-\mathbf{W}_{\mathbf{U}}\|$ $\leq\frac{2\|\widehat{\mathbf{X}}^{\top}\mathbf{X}-\mathbf{W}_{\mathbf{U}}\bm{\Lambda}\|}{\lambda_{i,\min}}\lesssim\frac{d^{1/2}(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})^{2}n_{i}}{q_{i}\lambda_{i,\min}^{2}}+\frac{d(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})\log^{1/2}n_{i}}{q_{i}^{1/2}\lambda_{i,\min}}$
@@
E.2 Proof of Lemma D.1
We first state a lemma for bounding the error of $\widehat{\mathbf{U}}^{(i)}$ as an estimate of $\mathbf{U}^{(i)}$; see Section E.3 for a proof.
Lemma E.4.
Consider the setting of Lemma D.1. Define $\mathbf{W}^{(i)}_{\mathbf{U}}=\underset{\mathbf{O}\in\mathcal{O}_{d}}{\arg\min}\|\widehat{\mathbf{U}}^{(i)}\mathbf{O}-\mathbf{U}^{(i)}\|_{F}$. We then have
$\widehat{\mathbf{U}}^{(i)}\mathbf{W}^{(i)}_{\mathbf{U}}-\mathbf{U}^{(i)}=\mathbf{E}^{(i)}\mathbf{U}^{(i)}(\bm{\Lambda}^{(i)})^{-1}+\mathbf{R}_{\mathbf{U}}^{(i)},$ (E.13)
@@
$\mathbf{X}_{\mathcal{U}_{i}}=\mathbf{U}_{\mathcal{U}_{i}}\bm{\Lambda}^{1/2},\quad\mathbf{X}^{(i)}=\mathbf{U}^{(i)}(\bm{\Lambda}^{(i)})^{1/2}\text{ and }\widehat{\mathbf{X}}^{(i)}=\widehat{\mathbf{U}}^{(i)}(\widehat{\bm{\Lambda}}^{(i)})^{1/2}.$
As $\mathbf{X}_{\mathcal{U}_{i}}\mathbf{X}_{\mathcal{U}_{i}}^{\top}=\mathbf{P}^{(i)}=\mathbf{X}^{(i)}\mathbf{X}^{(i)\top}$, there exists an orthogonal $\widetilde{\mathbf{W}}^{(i)}$ such that $\mathbf{X}_{\mathcal{U}_{i}}=\mathbf{X}^{(i)}\widetilde{\mathbf{W}}^{(i)}$. Define
$\breve{\mathbf{W}}^{(i)}=\underset{\mathbf{O}\in\mathcal{O}_{d}}{\arg\min}\|\widehat{\mathbf{X}}^{(i)}\mathbf{O}-\mathbf{X}^{(i)}\|_{F},$ (E.14)
@@
$\mathbf{W}^{(i)}=\underset{\mathbf{O}\in\mathcal{O}_{d}}{\arg\min}\|\widehat{\mathbf{X}}^{(i)}\mathbf{O}-\mathbf{X}_{\mathcal{U}_{i}}\|_{F}\text{ and }\mathbf{W}^{(i)}_{\mathbf{U}}=\underset{\mathbf{O}\in\mathcal{O}_{d}}{\arg\min}\|\widehat{\mathbf{U}}^{(i)}\mathbf{O}-\mathbf{U}^{(i)}\|_{F}.$
Note that $\mathbf{W}^{(i)}=\breve{\mathbf{W}}^{(i)}\widetilde{\mathbf{W}}^{(i)}$. Next, by Eq. (E.13) we have
$\widehat{\mathbf{X}}^{(i)}\mathbf{W}^{(i)}-\mathbf{X}_{\mathcal{U}_{i}}=$ $[\widehat{\mathbf{X}}^{(i)}\breve{\mathbf{W}}^{(i)}-\mathbf{X}^{(i)}]\widetilde{\mathbf{W}}^{(i)}$
$=$ $\Big{[}[(\widehat{\mathbf{U}}^{(i)}\mathbf{W}_{\mathbf{U}}^{(i)}-\mathbf{U}^{(i)})+\mathbf{U}^{(i)}](\mathbf{W}_{\mathbf{U}}^{(i)})^{\top}(\widehat{\bm{\Lambda}}^{(i)})^{1/2}(\breve{\mathbf{W}}^{(i)}-\mathbf{W}^{(i)}_{\mathbf{U}})$
$+[(\widehat{\mathbf{U}}^{(i)}\mathbf{W}_{\mathbf{U}}^{(i)}-\mathbf{U}^{(i)})+\mathbf{U}^{(i)}](\mathbf{W}_{\mathbf{U}}^{(i)})^{\top}[(\widehat{\bm{\Lambda}}^{(i)})^{1/2}\mathbf{W}^{(i)}_{\mathbf{U}}-\mathbf{W}^{(i)}_{\mathbf{U}}(\bm{\Lambda}^{(i)})^{1/2}]$
$+(\widehat{\mathbf{U}}^{(i)}\mathbf{W}^{(i)}_{\mathbf{U}}-\mathbf{U}^{(i)})(\bm{\Lambda}^{(i)})^{1/2}\Big{]}\widetilde{\mathbf{W}}^{(i)}$
$=$ $\mathbf{E}^{(i)}\mathbf{U}^{(i)}(\bm{\Lambda}^{(i)})^{-1/2}\widetilde{\mathbf{W}}^{(i)}+\mathbf{R}^{(i)}$
$=$ $\mathbf{E}^{(i)}\mathbf{X}^{(i)}(\mathbf{X}^{(i)\top}\mathbf{X}^{(i)})^{-1}\widetilde{\mathbf{W}}^{(i)}+\mathbf{R}^{(i)}$
@@
$\mathbf{R}^{(i)}=$ $\Big{[}\mathbf{R}^{(i)}_{\mathbf{U}}(\bm{\Lambda}^{(i)})^{1/2}+[(\widehat{\mathbf{U}}^{(i)}\mathbf{W}_{\mathbf{U}}^{(i)}-\mathbf{U}^{(i)})+\mathbf{U}^{(i)}](\mathbf{W}_{\mathbf{U}}^{(i)})^{\top}(\widehat{\bm{\Lambda}}^{(i)})^{1/2}(\breve{\mathbf{W}}^{(i)}-\mathbf{W}^{(i)}_{\mathbf{U}})$
$+[(\widehat{\mathbf{U}}^{(i)}\mathbf{W}_{\mathbf{U}}^{(i)}-\mathbf{U}^{(i)})+\mathbf{U}^{(i)}](\mathbf{W}_{\mathbf{U}}^{(i)})^{\top}[(\widehat{\bm{\Lambda}}^{(i)})^{1/2}\mathbf{W}^{(i)}_{\mathbf{U}}-\mathbf{W}^{(i)}_{\mathbf{U}}(\bm{\Lambda}^{(i)})^{1/2}]\Big{]}\widetilde{\mathbf{W}}^{(i)}.$
By Lemma E.4, Lemma E.2 and Lemma E.3 we obtain
$\|\mathbf{R}^{(i)}\|$ $\lesssim\|\mathbf{R}^{(i)}_{\mathbf{U}}\|\cdot\|\bm{\Lambda}^{(i)}\|^{1/2}+[\|\widehat{\mathbf{U}}^{(i)}\mathbf{W}_{\mathbf{U}}^{(i)}-\mathbf{U}^{(i)}\|+\|\mathbf{U}^{(i)}\|]\cdot\|\widehat{\bm{\Lambda}}^{(i)}\|^{1/2}\cdot\|\breve{\mathbf{W}}^{(i)}-\mathbf{W}^{(i)}_{\mathbf{U}}\|$
$+[\|\widehat{\mathbf{U}}^{(i)}\mathbf{W}_{\mathbf{U}}^{(i)}-\mathbf{U}^{(i)}\|+\|\mathbf{U}^{(i)}\|]\cdot\|(\widehat{\bm{\Lambda}}^{(i)})^{1/2}\mathbf{W}^{(i)}_{\mathbf{U}}-\mathbf{W}^{(i)}_{\mathbf{U}}(\bm{\Lambda}^{(i)})^{1/2}\|$
@@
$\|\mathbf{R}^{(i)}\|_{2\to\infty}$ $\lesssim\|\mathbf{R}^{(i)}_{\mathbf{U}}\|_{2\to\infty}\cdot\|\bm{\Lambda}^{(i)}\|^{1/2}+[\|\widehat{\mathbf{U}}^{(i)}\mathbf{W}_{\mathbf{U}}^{(i)}-\mathbf{U}^{(i)}\|_{2\to\infty}+\|\mathbf{U}^{(i)}\|_{2\to\infty}]\cdot\|\widehat{\bm{\Lambda}}^{(i)}\|^{1/2}\cdot\|\breve{\mathbf{W}}^{(i)}-\mathbf{W}^{(i)}_{\mathbf{U}}\|$
$+[\|\widehat{\mathbf{U}}^{(i)}\mathbf{W}_{\mathbf{U}}^{(i)}-\mathbf{U}^{(i)}\|_{2\to\infty}+\|\mathbf{U}^{(i)}\|_{2\to\infty}]\cdot\|(\widehat{\bm{\Lambda}}^{(i)})^{1/2}\mathbf{W}^{(i)}_{\mathbf{U}}-\mathbf{W}^{(i)}_{\mathbf{U}}(\bm{\Lambda}^{(i)})^{1/2}\|$
@@
$\widehat{\mathbf{U}}$ $=\mathbf{A}\widehat{\mathbf{U}}\widehat{\bm{\Lambda}}^{-1}=(\mathbf{U}\bm{\Lambda}\mathbf{U}^{\top}+\mathbf{E})\widehat{\mathbf{U}}\widehat{\bm{\Lambda}}^{-1}=\mathbf{U}\mathbf{U}^{\top}\widehat{\mathbf{U}}+\mathbf{U}(\bm{\Lambda}\mathbf{U}^{\top}\widehat{\mathbf{U}}-\mathbf{U}^{\top}\widehat{\mathbf{U}}\widehat{\bm{\Lambda}})\widehat{\bm{\Lambda}}^{-1}+\mathbf{E}\widehat{\mathbf{U}}\widehat{\bm{\Lambda}}^{-1}.$